A particle P moves on the curve with polar equation r = 1/ (2 - sinx) . Given that at any instant t, during the motion, r^2 (dx/dt) = 4,
(i) write an expression for r(dx/dt) in terms of x.
(ii) Show that dr/dt = 4cosx and 1/3 <=r<=1.
(iii) Find the speed of P when x = 0.
(iv) Prove that the force acting on P is directed towards the pole.
1
Expert's answer
2021-01-21T00:05:48-0500

"r=\frac{1}{2-\sin{x}}" (1)

"r^2\frac{dx}{dt}=4" (2)


(i)

"r\frac{dx}{dt}=\frac{4}{r}=4(2-\sin{x})"


(ii)

"\frac{dr}{dt}=\frac{dr}{dx}\frac{dx}{dt}="

"=\frac{(-1)(-\cos{x})}{(2-\sin{x})^2}*4(2-\sin{x})^2=4\cos{x}"


"-1\leqslant-\sin{x}\leqslant1"

"1\leqslant2-\sin{x}\leqslant3"

"1\geqslant\frac{1}{2-\sin{x}}\geqslant\frac{1}{3}"

"\frac{1}{3}\leqslant r \leqslant 1"


(iii)

"v - speed(velocity)"

"v=\sqrt{v_r^2+v_x^2}=\sqrt{(\frac{dr}{dt})^2+(r\frac{dx}{dt})^2}="

"=\sqrt{16\cos^2{x}+16(2-\sin{x})^2}="

"=\sqrt{16+16*4-64\sin{x}}=\sqrt{80-64\sin{x}}"

"v(x=0)=\sqrt{80}=4\sqrt{5}"


(iv)

"a- acceleration"

"a_x=r(\frac{d^2x}{dt^2})^2+2\frac{dr}{dt}\frac{dx}{dt}="

"=\frac{1}{2-\sin{x}}\frac{d}{dt}(4(2-\sin{x})^2)+2*4\cos{x}*4(2-\sin{x})^2="

"=\frac{4*2(2-\sin{x})}{2-\sin{x}}(-\cos{x})\frac{dx}{dt}+32\cos{x}(2-\sin{x})^2="

"=(-8)\cos{x}*4(2-\sin{x})^2+32\cos{x}(2-\sin{x})^2=0"


"a_r=\frac{d^2r}{dt^2}-r(\frac{dx}{dt})^2="

"=\frac{d}{dt}(4\cos{x})-\frac{1}{2-\sin{x}}(4(2-\sin{x})^2)^2="

"=-4\sin{x}\frac{dx}{dt}-16(2-\sin{x})^3="

"=-4\sin{x}*4(2-\sin{x})^2-16(2-\sin{x})^3="

"=-16(2-\sin{x})^2(\sin{x}+2-\sin{x})="

"=-32(2-\sin{x})^2"


"F=ma" , so the force acting on P is directed towards the pole